### v2

# arXiv:2512.20043v2[cs.AI] 01 Mar 2026

Abstract
Symmetry is fundamental to understanding physical systems and can improve performance and sample efficiency in machine learning. Both pursuits require knowledge of the underlying symmetries in data, yet discovering these symmetries automatically is challenging. We propose LieFlow, a novel framework that reframes symmetry discovery as a distribution learning problem on Lie groups. Instead of searching for the symmetry generators, our approach operates directly in group space, modeling a symmetry distribution over a large hypothesis group $G$. The support of the learned distribution reveals the underlying symmetry group $H\subseteq G$. Unlike previous works, LieFlow can discover both continuous and discrete symmetries within a unified framework, without assuming a fixed Lie algebra basis or a specific distribution over the group elements. Experiments on synthetic 2D and 3D point clouds and ModelNet10 show that LieFlow accurately discovers continuous and discrete subgroups, significantly outperforming a state-of-the-art baseline, LieGAN, in identifying discrete symmetries.
Machine Learning, ICML
1 Introduction
Symmetry has long been central to mathematics and physics and, more recently, has featured prominently within machine learning. Classically, to model physical systems, the underlying symmetry is first identified and categorized; Noether’s theorem then (Noether, 1918) establishes the existence of corresponding conservation laws, which can be used to understand and predict the system. In machine learning, symmetries serve as powerful inductive biases in equivariant neural networks (Cohen and Welling, 2016; Kondor and Trivedi, 2018; Weiler and Cesa, 2019). These architectures leverage known symmetries in the data, such as rotational invariance in molecular structures (Thomas et al., 2018; Satorras et al., 2021) or translational equivariance in images (LeCun et al., 1998), to achieve superior performance with fewer parameters and training samples (Bronstein et al., 2021).
Figure 1: LieFlow discovers symmetries from data by assuming a hypothesis Lie group $G$ and learning the distribution of transformations over $G$ via flow matching over the Lie group, conditioned on data samples $x$.
A critical limitation in both physics and machine learning is that the exact symmetry group must be known a priori. In practice, the underlying symmetries are often unknown, approximate, or domain-specific, e.g. hidden symmetries in physical systems (Gross, 1996; Liu and Tegmark, 2022), non-obvious chemical properties (Davies et al., 2016), and partial symmetries in 3D geometry (Mitra et al., 2006, 2013). Many applications can thus benefit from the ability to automatically discover and characterize symmetries from data, without relying on domain expertise or manual specification.
Several prior works have focused on symmetry discovery in restricted settings, such as roto-translations in images (Rao and Ruderman, 1998; Miao and Rao, 2007), commutative Lie groups (Cohen and Welling, 2014), or finite groups (Zhou et al., 2020; Karjol et al., 2024). Recent studies on learning Lie groups also have some limitations: Benton et al. (2020) assume a fixed Lie algebra basis and a uniform distribution over the coefficients; Dehmamy et al. (2021) produce non-interpretable symmetries; Yang et al. (2023) assume a Gaussian distribution over the Lie algebra coefficients which is largely limited to continuous groups and use potentially unstable adversarial training; and Allingham et al. (2024) use a fixed Lie algebra basis and evaluate only over images.
In this work, we propose LieFlow, a new framework that fundamentally reframes symmetry discovery as a distribution learning problem on Lie groups. Conceptually, this shifts the paradigm of estimating generators to recovering group structure via support learning in group space. Specifically, we learn a distribution over a larger symmetry group, which serves as a hypothesis space of symmetries. The support of this distribution corresponds to the actual transformations observed in the data. We achieve this via flow matching directly on Lie groups (Figure 1).
Unlike standard flow matching (Lipman et al., 2023), which operates in data space, LieFlow learns a flow on a Lie group $G$ which is acting on the data space, conditioned on a data point, enabling the generation of new plausible data that respect the underlying symmetry structure. Thanks to the flexibility of flow matching, LieFlow can capture both continuous and discrete symmetries within a unified framework and accurately model the multi-modal nature of distributions supported on subgroups. We observe that discrete symmetry groups can be particularly challenging since cancellation among the different flow vectors results in small velocity until late in the flow, creating a “last-minute convergence” phenomenon. We counter this issue by proposing a novel time schedule for flow matching.
Qualitative and quantitative experimental results on synthetic 2D and 3D datasets and real-world dataset ModelNet10 (Wu et al., 2015) show that LieFlow can discover discrete and continuous symmetries accurately, and achieve significant improvements over LieGAN (Yang et al., 2023) baseline in identifying discrete symmetries.
Our key contributions are as follows:
  • We propose a novel formulation of symmetry discovery as a flow matching problem on Lie groups. Our formulation covers continuous, discrete, approximate, and partially observed symmetries.
  • We identify a “last-minute convergence” phenomenon for discrete symmetries and propose a novel time scheduling method to address it.
  • Extensive experiments on synthetic and real-world datasets demonstrate that LieFlow can accurately discover underlying symmetry groups, and achieve significant improvements over LieGAN (Yang et al., 2023) baseline in identifying discrete symmetries.
2 Related Work
Symmetry Discovery.
Many works on symmetry discovery differ on the types of symmetries that they can learn. Early works (Rao and Ruderman, 1998; Miao and Rao, 2007) used sequences of transformed images to learn Lie groups in an unsupervised way, but were limited to 2D roto-translations. Some recent works (Zhou et al., 2020; Karjol et al., 2024) can only discover finite groups, while others learn partially observed symmetries, i.e. subsets of known groups (Benton et al., 2020; Romero and Lohit, 2022; Chatzipantazis and Pertigkiozoglou, 2023). Shaw et al. (2024) discover continuous and discrete symmetries of machine learning functions beyond affine transformations, but focus on Killing type generators rather than probability flows on a Lie group. In contrast, LieFlow is flexible enough to learn both continuous and discrete symmetries, highly multi-modal distributions, and also partially observed symmetries via flow matching.
Several related works focus on continuous Lie groups but suffer from other limitations. Dehmamy et al. (2021) use Lie algebras in CNNs but learn the symmetries end-to-end along with the task function, leading to non-interpretable symmetries. Moskalev et al. (2022) extract Lie group generators from a neural network to analyze the learned equivariance, rather than discovering which symmetries exist in the dataset. Otto et al. (2023) discover symmetries of learned models via linear algebraic operators. Ko et al. (2024) learn infinitesimal generators of continuous symmetries from data via Neural ODEs and a task-specific validity score, capturing non-affine and approximate symmetries. Unlike these approaches, LieFlow models a full distribution over transformations within a prescribed matrix Lie group.
Yang et al. (2023) and Allingham et al. (2024) are most related to our work as they also learn distributions over transformations. Yang et al. (2023) propose LieGAN, which uses adversarial training to learn the Lie algebra basis while assuming a fixed distribution over the coefficients, typically Gaussian. This makes LieGAN well-suited for continuous symmetries; while it can discover discrete groups, doing so requires a carefully chosen prior distribution such as a multi-modal delta mixture. Allingham et al. (2024) first learn a canonicalization function to obtain a prototype as the orbit representative, then learn a generative model over transformations conditioned on the orbit representative. LieFlow, on the other hand, learns a marginal transformation distribution across all orbits directly via flow matching, without assuming a fixed distribution or learning explicit prototypes.
Generative Models on Manifolds.
Several recent works have extended continuous normalizing flows to manifolds (Gemici et al., 2016; Mathieu and Nickel, 2020; Falorsi, 2021) but rely on computationally expensive likelihood-based training. Some works (Rozen et al., 2021; Ben-Hamu et al., 2022; Chen and Lipman, 2024) have proposed simulation-free training methods; in particular, Chen and Lipman (2024) considers flow matching on general Riemannian manifolds. However, none of these works specifically consider Lie groups or the task of discovering symmetries from data. Other flow matching works directly incorporate known symmetries (Klein et al., 2023; Song et al., 2023; Bose et al., 2023) and do not infer symmetry from data. Another class of methods considers score matching and flow matching directly on Lie groups as data spaces (Zhu et al., 2025; Bertolini et al., 2025; Sherry and Smets, 2025). While methodologically similar to our work, they assume prior knowledge of the symmetry group and use specialized implementations for different Lie groups.
3 Preliminaries
Lie Group.
A Lie group is a group that is also a smooth manifold, such that group multiplication and inversion are smooth maps. For example, the group of planar rotations $\mathrm{SO}(2)$ is a Lie group parameterized by angle $\theta$, with elements $R_{\theta}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}$. Each Lie group $G$ has an associated Lie algebra $\mathfrak{g}$, which is the tangent space at the identity. The Lie algebra captures the local (infinitesimal) group structure.
The exponential map $\exp\colon\mathfrak{g}\to G$ maps the Lie algebra to the Lie group. For matrix Lie groups, which we consider exclusively in this work, the exponential map is the matrix exponential $\exp(A)=\sum_{k=0}^{\infty}\frac{A^{k}}{k!}$. If the exponential map is surjective, we can restrict its domain and define the logarithm map $\log\colon G\supset U\to\mathfrak{g}$ for the neighborhood $U$ around the identity such that $\exp(\log(g))=g$ for $g\in U$. Note that any Lie algebra element $A$ can be written as a linear combination of the Lie algebra basis. This allows us to generate the entire connected component of the identity in the Lie group using the Lie algebra, and thus the Lie algebra basis $L_{i}\in\mathfrak{g}$ is also called the (infinitesimal) generators of the Lie group.
Flow Matching.
[…]
$x_{t}=(1-t)x_{0}+tx_{1},\quad\dot{x}_{t}=x_{1}-x_{0},$ (1)
defining a stochastic process $\{x_{t}\}$, also referred to as a probability path.
In Flow Matching, we define the velocity field $u_{t}$ that generates the desired probability path, and train a neural velocity field $v_{t}^{\theta}$ to match it:
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t,x}\,\|v_{t}^{\theta}(x)-u_{t}(x)\|^{2}.$ (2)
Since $u_{t}$ can be written as the conditional expectation of these per-sample derivatives, $u_{t}(x)=\mathbb{E}[\dot{x}_{t}\mid x_{t}=x],$ we obtain the practical Conditional Flow Matching (CFM) loss
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,x_{0},x_{1}}\,\|v_{t}^{\theta}(x_{t})-\dot{x}_{t}\|^{2},$ (3)
where $x_{t}$ and $\dot{x}_{t}$ come from the chosen interpolation.
A key advantage of FM is that the transport dynamics can be learned entirely in a self-supervised manner, since interpolations between noise and data provide training targets without requiring access to likelihoods or score functions.
4 Symmetry Discovery via Flow Matching
4.1 Motivation and Formalism
A central challenge in symmetry discovery is not merely to estimate transformation parameters, but to determine which transformations are symmetries at all. In realistic settings, we are not given paired observations of original and transformed data, and the space of candidate transformations is typically much larger than the true symmetry group expressed in the data. We address this challenge by reframing symmetry discovery as a distribution learning problem on groups. Rather than directly estimating generators or invariants, we learn a probability distribution over a large hypothesis group $G$, whose support concentrates on the true symmetry subgroup $H\subseteq G$. We realize this idea using flow matching on Lie groups, which provides a flexible and self-supervised mechanism for learning which transformations preserve the data distribution.
Problem Formulation.
Our goal is to recover the unknown group of transformations $H$ that preserves the data distribution $q$ on data space $\mathcal{X}$. We first assume a large hypothesis group $G$ acting on $\mathcal{X}$ that contains hypothetical symmetries. We wish to find the subgroup $H\subseteq G$ such that $q(hx)=q(x)$ for all $x\in\mathcal{X}$ and $h\in H$.
Discovering Symmetry as Distribution Learning.
We frame the problem of finding $H$ as learning a distribution over the hypothesis group $G$ that concentrates on the subgroup $H$. To achieve this, we train an FM model that maps a chosen prior distribution over $G$ to a distribution $p_{\theta}(g)$ supported on $H$. The structure of the discovered symmetries emerges from the concentration pattern of $p_{\theta}(g)$: continuous symmetries yield a distribution spreading across the manifold of valid transformations, while discrete subgroups exhibit sharp peaks at a finite set of group elements (e.g. the four-fold rotations of $C_{4}$). Thus, LieFlow effectively filters the large hypothesis group down to the subgroup $H\subseteq G$ consistent with the data.
Hypothesis Group and Prior over Symmetries.
We assume that the hypothesis group $G$ is connected, ensuring a well-defined logarithm map for compact groups. For the prior distribution $p$ over the group $G$, we use a uniform prior for compact groups and bounded priors over the Lie algebra coefficients for non-compact groups. When the exponential map is not surjective, we define the prior implicitly by sampling from the Lie algebra $\mathfrak{g}$ and applying the exponential map to construct group elements, ensuring well-defined training targets. In experiments, we consider $G=\mathrm{SO}(2)$ (planar rotations), $\mathrm{GL}(2)$ (invertible matrices) for the 2D datasets and $G=\mathrm{SO}(3)$ (3D rotations) for the 3D datasets.
Interpolation Paths Along Orbits.
We now describe how to construct interpolation paths that respect the group structure. Consider a data sample $x_{1}\sim q$ and a transformation $g\sim p(G)$ drawn from the prior distribution. We apply the transformation to obtain $x_{0}=gx_{1}$, which serves as the starting point of our flow. We wish to learn to flow from $x_{0}$ to $x_{1}$ using only transformations in $G$. Let $A=\log(g^{-1})\in\mathfrak{g}$ be the Lie algebra element that generates the inverse transformation. We define the interpolation in the data space via the group action:
$x_{t}=\exp(tA)x_{0},\quad t\in[0,1],$ (4)
At $t=0$, we have $x_{0}=gx_{1}$, and at $t=1$, we recover $x_{1}$ since $\exp(A)=g^{-1}$. Marginalizing over $g$, this defines a conditional $p_{t}(x_{t}\mid x_{1})$ that stays on the group orbit of $x_{1}$.
The Target Velocity Field.
The derivative $\frac{d}{dt}x_{t}\in T_{x_{t}}\mathcal{X}$ defines a vector field along the path $\{x_{t}\}_{t\in[0,1]}$. Unlike standard flow matching, where the network outputs a velocity in data space, we configure our network to output a Lie algebra element $A\in\mathfrak{g}$ directly. That is, we set the conditional target field to $u_{t}(x_{t}\ |x_{1})=A$.
This choice has two advantages. First, the update rule $x_{t+\Delta_{t}}=\exp(\Delta_{t}A_{t})x_{t}$ ensures that the trajectory stays on the group orbit of $x_{1}$. Second, we can accumulate the transformations to obtain group elements in $H$ (Line 9, Algorithm 2). Note that $A$ is constant along the path, so the network learns to predict the Lie algebra element generating the transformation from $x_{0}$ to $x_{1}$, regardless of $t$. The velocity field $\frac{d}{dt}x_{t}$ is obtained from $A$ via the differential of the group action; see Appendix A for more details.
Objective.
The CFM objective on Lie groups is similar to the Euclidean case and is given by
$\mathcal{L}_{\mathrm{LieCFM}}=\mathbb{E}_{t,q,p_{t}(x_{t}\mid x_{1})}\|v_{t}^{\theta}(x_{t})-u_{t}(x_{t}\mid x_{1})\|^{2}_{\mathcal{G}},$ (5)
where $x_{1}\sim q$, $x_{t}$ is defined by Eq. (4), $p_{t}(x_{t}\mid x_{1})$ is the distribution of points along the paths, and $\mathcal{G}$ is a Riemannian metric on the Lie group, which equips the manifold with an inner product on each tangent space, allowing us to measure distances along the manifold. We use a left-invariant metric as it is completely determined by the inner product on the Lie algebra $\mathfrak{g}$ and the push-forward of the left action is an isometry, making computations simpler. The specific form of this metric is determined by the structure of the Lie group under consideration.
Although transformations are sampled from a broad hypothesis group during training, the flow-matching objective implicitly filters out transformations that are not true symmetries of the data. When a sampled transformation preserves the data distribution, applying it to a data point produces another statistically valid sample, and the resulting interpolation paths induce consistent training signals across many data points. In contrast, transformations that do not preserve the data distribution produce inconsistent targets: different data points yield incompatible directions, which cancel out when the model averages training signals. As a result, only transformations that are globally consistent with the data admit a coherent vector field, and the learned symmetry distribution concentrates its probability mass on the true symmetry group.
4.2 Training and Generation with LieFlow
Algorithm 1 Training
1:repeat
2:  $x_{1}\sim q$
3:  $g\sim p(G)$
4:  $t\in[0,1)$
5:  $x_{0}=gx_{1}$
6:  $A=\log(g^{-1})$
7:  $x_{t}=\exp(tA)x_{0}$
8:  Take gradient descent step on $\nabla_{\theta}\left\|v_{\theta}(x_{t},t)-A\right\|^{2}$
9:until converged
Training.
Using the interpolation paths and the Lie group flow matching objective, we can now derive the training and sampling algorithms for our method. The training algorithm (Algorithm 1) directly implements the objective by sampling $x_{1}$ from the data distribution, sampling a transformation $g\sim p(G)$, and constructing $x_{0}=gx_{1}$. The target Lie algebra group element $A$ can be computed analytically using the sampled group element $g$. We sample a timestep $t\in[0,1)$ and construct a point $x_{t}=\exp(tA)x_{0}$ on the curve. Given $x_{t}$ and $t$, the model predicts $A$. We later show a modified time sampling procedure that works well for discovering discrete subgroups in Section 5.4.
Sampling and Generation.
During sampling (Algorithm 2), we integrate over the learned vector field to produce a new sample $x_{1}^{\prime}$. Starting from $x_{0}=gx_{1}$ for a random $g\sim p(G)$, we iteratively apply the predicted Lie algebra elements using Euler integration with the exponential map $x_{t+\Delta t}^{\prime}=\exp(\Delta_{t}A_{t})x_{t}^{\prime}$. This ensures that the trajectory remains on the group orbit.
Algorithm 2 Generating data
[…]
4:$M=I$
[…]
6:for $t=\{0,\Delta_{t},2\Delta_{t},\dots,(T-1)\Delta_{t}\}$ do
7:  $A_{t}=v^{\theta}(x_{t}^{\prime},t)$
8:  $x_{t+\Delta_{t}}^{\prime}=\exp(\Delta_{t}A_{t})x_{t}^{\prime}$
9:  $M=\exp(\Delta_{t}A_{t})M$ $\triangleright$ Accumulate transforms
10:end for
11:return $x_{1}^{\prime},M$
Algorithm 3 Generating Group Elements
[…]
5:return MgMg
We also accumulate the product of all applied transformations into $M$ such that at the end of sampling, $x_{1}^{\prime}=Mx_{0}^{\prime}$. The transformation $M$ is used for generating new group elements (Algorithm 3).
Algorithm 3 shows how to generate group elements $h$ consistent with the target group $H$. Given a data sample $x_{1}$, we transform it by $g\sim p(G)$ to obtain $x_{0}=gx_{1}$. We then run Alg. 2 partially (lines 3-10) on $x_{0}$, obtaining outputs $x_{1}^{\prime}$ and $M$. The key observation is that the composed transform MgMg forms a group element $h\in H$ that maps $x_{1}$ to $x_{1}^{\prime}$.
4.3 Handling Partially Observed Symmetries
The main assumption of this paper is that there exists a group $H$ to which the data distribution is invariant. This requires that the data be spread uniformly along the orbits of an entire group. In practice, however, data distributions can exhibit partial symmetries, i.e. $q(hx)=q(x)$ holds for only some $x\in\mathcal{X}$ and $h\in H$. We discuss how our method can be generalized to handle this case.
Formally, we consider the scenario where there exists a maximal subset $H_{x}\subseteq H$ for each $x\in\mathcal{X}$ such that for all $h\in H_{x}$, $q(hx)=q(x)$. For example, consider a data distribution of three arrow shapes: $q(\leftarrow)=q(\uparrow)=q(\rightarrow)=1/3$, and the group $H=\langle h\rangle$, where $h$ acts on data by clockwise $\pi/2$ rotation. Then we have $H_{\leftarrow}=\{e,h,h^{2}\}$, $H_{\uparrow}=\{e,h,h^{3}\}$, $H_{\rightarrow}=\{e,h^{2},h^{3}\}$. The goal of symmetry discovery then becomes, in addition to finding a subgroup $H\subset G$, identifying the input-dependent “symmetry subset” $H_{x}\subseteq H$ for any $x$.
Note that our method does not learn the explicit algebraic structure of $H$, but instead learns a distribution over the hypothesis group $G$ and identifies $H$ by (the closure of) the support of the learned distribution density. Analogously, to discover the partial symmetries stated above, it suffices to learn a conditional distribution over $G$, $p_{\theta}(\cdot|x)$, whose support at each $x$ reveals the symmetry subset $H_{x}$.
In Algorithm 3, we first sample $x_{1}$ from the data distribution $q$ and then sample $h=Mg$ that maps $x_{1}$ to another point $x_{1}^{\prime}$ with an equal density. In other words, we can view $h$ as a sample from a conditional distribution $p_{\theta}(\cdot|x_{1})$. In our default setting of exact global symmetry, we do not need to specify the condition $x_{1}$ because the same symmetry group applies at every point. Algorithm 3 then simple performs ancestral sampling from the marginal $p_{\theta}(\cdot)=\int_{\mathcal{X}}p_{\theta}(\cdot|x_{1})q(x_{1})dx_{1}$. In the setting with partial symmetries, the same procedure instead queries an input-dependent symmetry subset $H_{x_{1}}\subset H$, revealed by the support of $p_{\theta}(\cdot|x_{1})$ at a given $x_{1}$. In Section 5.3, we show that our method can still discover the support of the underlying symmetry group even when the data only demonstrates partial symmetries.
(a) $\mathrm{SO}(2)$ to $C_{4}$
(b) $\mathrm{GL}(2,\mathbb{R})^{+}$ to $C_{4}$ (Lie algebra coefficients based sampling)
(c) $\mathrm{GL}(2,\mathbb{R})^{+}$ to $C_{4}$ (matrix composition based sampling)
(d) $\mathrm{SO}(2)$ to partial $C_{4}$
Figure 2: Histograms of $5,000$ learned angles for 2D datasets, where the peaks correspond to the elements of the target symmetry group $C_{4}$. LieFlow successfully recovers $C_{4}$ symmetry group from (a) a compact hypothesis group $\mathrm{SO}(2)$ (b) a non-compact hypothesis group, (c) different sampling strategies, and (d) even when the full symmetry group is partially observed. The peaks correspond to the elements of the target symmetry group $C_{4}$.
5 Experiments
To evaluate LieFlow, we consider a variety of datasets consisting of 2D and 3D point clouds with different underlying ground truth symmetry groups. We also show our method can succeed with different hypothesis groups, in settings with partially observed symmetries, where only a subset of group transformations are observed in data, and in settings where observations are noisy.
5.1 Datasets
We evaluate LieFlow on both synthetic 2D and 3D point clouds and on ModelNet10 (Wu et al., 2015), which contains 3D models of real-world objects. For both the synthetic and real-world objects, we generate several datasets by transforming the canonical objects with group elements sampled from a target ground truth symmetry group $H$, see Figures 56 in Appendix B.
Synthetic datasets
We consider two datasets, each generated from a single canonical object: a 2D arrow and a 3D irregular tetrahedron with no self-symmetries. For the 2D arrow, we add data points given by $H=C_{4}$ transformations and perform flow matching with hypothesis groups $\mathrm{SO}(2)$ and $\mathrm{GL}(2,\mathbb{R})^{+}$. For $\mathrm{SO}(2)$, we uniformly sample angles in $0,2\pi)$ to generate group elements. For $\mathrm{GL}(2,\mathbb{R})^{+}$, which is non-compact, we consider two different prior distributions to sample group elements $g$: (1) Lie algebra coefficients based sampling (L): uniformly sampling the coefficients of the four orthogonal basis vectors and applying the exponential map; (2) Matrix composition based sampling (M): independently sampling rotation, scaling, and shear matrices and multiplying them. See Appendix C for more details. For the 3D dataset, we attempt to discover four different ground truth symmetry groups: tetrahedral $\mathrm{Tet}$, octahedral $\mathrm{Oct}$, icosahedral $\mathrm{Ico}$ and $SO(2)$. We use $\mathrm{SO}(3)$ as a hypothesis group.
ModelNet10
To demonstrate that LieFlow can handle complex real-world objects of diverse shapes, we consider ModelNet10 dataset(Wu et al., 2015). We first randomly select $10$ objects from each of the ten categories of ModelNet10 as canonical shapes, and downsample the point clouds to $128$ points. See Figure 6 in Appendix B for example samples. We consider five different target symmetry groups $\mathrm{Tet}$, $\mathrm{Oct}$, $\mathrm{Ico}$, the $\mathrm{SO}(2)$ subgroup stabilizing the $z$ axis and $\mathrm{SO}(2)$ stabilizing $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$.
5.2 Evaluation
In order to properly measure the quality of discovered symmetries, we propose using the Wasserstein-1 distance between the generated group elements (sampled via Algorithm 3) and the ground truth elements. Given two distributions $P$ and $Q$ over a metric space $(\mathcal{M},d_{G})$, let $\Pi(P,Q)$ be the set of all couplings between $P$ and $Q$, the Wasserstein-1 distance is defined as:
$W_{1}(P,Q)=\inf_{\gamma\in\Pi(P,Q)}\mathbb{E}_{(x,y)\sim\gamma}[d_{G}(x,y)],$ (6)
and it can be computed using the POT library in Python
For all experiments, we use a simple five-layer MLP with RELU activations. As a baseline, we compare against LieGAN (Yang et al., 2023), a recent state-of-the-art method for learning Lie group symmetries using a GAN. We set the number of generators in LieGAN to be the same dimension as our hypothesis group, and keep the architecture of the discriminator the same as our predictive network for a fair comparison. See Appendix C for more details.
5.3 Results on Synthetic 2D Datasets
Figures 2(a), 2(b), and 2(c) show the histograms of learned angles for the 2D dataset with target group $C_{4}$. All histograms correctly display four peaks at the $C_{4}$ group elements, confirming that LieFlow discovers the underlying symmetries across different hypothesis groups and prior distributions.
Table 1: Wasserstein-1 distance between generated group elements and ground truth group elements in 2D experiments (lower is better). We use 11 and $4$ LieGAN generators, corresponding to the dimension of $\mathrm{SO}(2)$ and $\mathrm{GL}(2,\mathbb{R})^{+}$, respectively. (L) and (M) indicate Lie algebra coefficients based sampling and matrix composition based sampling, respectively.
Method $\mathrm{SO}(2)\to C_{4}$ $\mathrm{GL}(2,\mathbb{R})^{+}\to C_{4}$(L) $\mathrm{GL}(2,\mathbb{R})^{+}\to C_{4}$(M)
LieGAN (1.gen) $1.688$ $1.688$ $1.688$
LieGAN (4.gen) $1.672$ $1.672$ $1.672$
LieFlow(Ours) $\mathbf{0.072}$ $\mathbf{1.321}$ $\mathbf{0.428}$
Table 1 shows the Wasserstein-1 distance between the generated group elements and the ground truth group elements. LieFlow outperforms LieGAN in all settings. We can see that using a smaller hypothesis group $\mathrm{SO}(2)$ leads to much better performance, matching our intuition that a more informative prior is beneficial. Even with the larger $\mathrm{GL}(2)$ hypothesis group, both sampling settings outperform LieGAN, although the matrix composition prior (M) is superior.
(a) $\mathrm{SO}(3)$ to $\mathrm{Tet}$
(b) $\mathrm{SO}(3)$ to $\mathrm{Oct}$
(c) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over z axis
(d) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$ axis
(e) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, without power schedule
(f) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule
Figure 3: 3D Datasets: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by Euler angles. The first two angles are represented spatially on the sphere using Mollweide projection and the color represents the third angle. The elements are canonicalized by the original random transformation and the ground truth elements of the target group are shown in circles with gray borders. The points are jittered with uniformly random noise to prevent overlapping. Our method recovers a variety of ground truth symmetry groups.
Partially Observed Symmetries. In this experiment, we consider the setting where only a subset of the transformations of the target group $C_{4}$ are present in the dataset, i.e. $C_{4}$ is only partially observed. Specifically, we construct the dataset by considering only $0,\frac{\pi}{2},\frac{3\pi}{2}$ degree rotations to the canonical arrow object with equal probability.
The results are shown in Figure 2(d). Even when only three transformations are observed in the dataset, LieFlow using Algorithm 3 can correctly recover all four elements of the group closure $C_{4}$.
5.4 Results on Synthetic 3D Datasets
Figures 3 visualize generated group elements for the different target symmetry groups on $\mathrm{SO}(3)$, where the points are generated from our flow model and the ground truth symmetries are shown as circles with gray borders. For the tetrahedral group (Figure 3(a)), octahedral group (Figure 3(b)), $\mathrm{SO}(2)$ group around the $z$-axis (Figure 3(c)), and $\mathrm{SO}(2)$ group around $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$ (Figure 3(d)), we can see that our model learns to correctly output transformations very close to the target group elements, similar to the 2D case. We find that training with a uniform time sampling procedure in Algorithm 1 fails to recover the icosahedral group and produces almost uniformly distributed elements over $\mathrm{SO}(3)$. We provide an improved sampling method in the next section.
(a) $\mathrm{SO}(3)$ to $\mathrm{Tet}$
(b) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$
Figure 4: Time progression of $x_{t}$ when generating $5$ samples over $10$ steps to illustrate the inference process. The object starts moving at later time steps in the discrete group case ((a)), whereas it gradually moves at each time step in the continuous group case ((b)). The gray object shows the original $x_{1}$.
Table 2(a) presents quantitative results, where LieFlow significantly outperforms LieGAN in all discrete group while achieving comparable performance on continuous groups.
Improved Time Sampling. To further analyze the failure of uniform time sampling, we first plot the time progression of transformed objects in Figure 4(a) and Figure 4(b). For discrete groups, the model only begins moving objects towards the group elements when $t$ is close to $1$, whereas for continuous groups, the transformation occurs gradually throughout the flow. We find that the entropy of the posterior remains stationary until $t$ approaches $1$ (see Appendix D for more details). This phenomenon of “last-minute convergence” makes learning discrete groups particularly challenging as the model needs to learn to produce a near-zero vector field for most time steps and only moves towards the correct mode at the end of inference.
Table 2: Wasserstein-1 distance between generated group elements and ground truth group elements in irregular tetrahedron (Section 5.4) and rotated ModelNet10 (Section 5.5) experiments (lower is better). $\mathrm{SO}(3)\to\mathrm{SO}(2)$(a) corresponds to rotations around $z$ axis and $\mathrm{SO}(3)\to\mathrm{SO}(2)$(b) corresponds to rotations around $(0,1/2,-\sqrt{3}/2)$ axis. $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/o pow.) indicates results without power time schedule and $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/ pow.) indicates results with power time schedule.
(a) 3D irregular tetrahedron (Section 5.4)
Method $\mathrm{SO}(3)\to\mathrm{Tet}$ $\mathrm{SO}(3)\to\mathrm{Oct}$ $\mathrm{SO}(3)\to\mathrm{SO}(2)$(a) $\mathrm{SO}(3)\to\mathrm{SO}(2)$(b) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/o pow.) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/ pow.)
LieGAN $2.326$ $2.198$ $0.0693$ $0.0309$ $1.861$ $1.861$
LieFlow(Ours) $\mathbf{0.0885}$ $\mathbf{0.0962}$ $\mathbf{0.0429}$ $\mathbf{0.0299}$ $\mathbf{0.319}$ $\mathbf{0.118}$
(b) ModelNet10 (Section 5.5)
Method $\mathrm{SO}(3)\to\mathrm{Tet}$ $\mathrm{SO}(3)\to\mathrm{Oct}$ $\mathrm{SO}(3)\to\mathrm{SO}(2)$(a) $\mathrm{SO}(3)\to\mathrm{SO}(2)$(b) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/o pow.) $\mathrm{SO}(3)\to\mathrm{Ico}$ (w/ pow.)
LieGAN $2.291$ $2.233$ $\mathbf{0.0291}$ $0.0621$ $1.828$ $1.828$
LieFlow(Ours) $\mathbf{0.113}$ $\mathbf{0.125}$ $0.0461$ $\mathbf{0.0472}$ $\mathbf{0.717}$ $\mathbf{0.156}$
Since most of the learning signal for discrete groups occurs near $t=1$, we leverage a time scheduling scheme to sample more timesteps near $t=1$. Instead of sampling from $\mathcal{U}(0,1)$, we use a power distribution with density $p(x)=nx^{n-1},x\sim\mathcal{U}(0,1)$, where $n$ controls the skewness (see Figure 13 in Appendix E). We find empirically that $n=5$ works best; see Appendix E.1. Figure 14 in Appendix E confirms that with the power time schedule, objects begin moving earlier in the flow. With this power schedule, our model correctly learns the icosahedral group as shown in Figure 3(f).
5.5 Results on ModelNet10
Table 2(b) shows the quantitative results for different target symmetry groups of real-world ModelNet10 objects, with qualitative visualizations in Appendix F.3. As with the synthetic 3D datasets, LieFlow significantly outperforms LieGAN on all discrete groups and achieves comparable performance for continuous groups, demonstrating its effectiveness in discovering symmetries in real-world datasets. Again we find that using the power time schedule over uniform sampling improves performance for the higher-order discrete icosahedral group.
Incomplete Observations
To evaluate the robustness of LieFlow to incomplete observations, we randomly mask a portion of the points in ModelNet10 point clouds. Table 3 shows the results for the challenging setting of discovering $\mathrm{Ico}$ with hypothesis group $\mathrm{SO}(3)$ with different masking ratios (0%, 5%, 10%). Even with $10\%$ masking, LieFlow still discovers the underlying symmetries with only moderate degradation in performance. See Figures 22(g)22(h) in the Appendix F.3 for visualizations.
Table 3: Incomplete observations: Wasserstein-1 distance between generated group elements and ground truth group elements for $\mathrm{SO}(3)\to\mathrm{Ico}$ with different masking ratios (lower is better).
Method 0% 5% 10%
LieFlow 0.156 0.186 0.212
6 Conclusion and Future Work
Our work demonstrates that expressive generative models like flow matching offer a promising avenue for symmetry discovery. We introduced LieFlow, a novel framework that formulates symmetry discovery as learning a distribution on Lie groups using flow matching. By learning a flow from a large hypothesis group $G$ to a distribution whose support matches the true symmetry subgroup $H\subseteq G$, our approach discovers both continuous and discrete symmetries within a unified framework, without making assumptions on the Lie algebra basis or the distribution of its coefficients.
Our experiments demonstrate that LieFlow successfully discovers a variety of symmetry groups across different settings, such as $C_{4}$ and partial $C_{4}$ symmetries in 2D datasets, and several discrete and continuous $\mathrm{SO(3)}$ subgroups acting on 3D space. In addition, our method also demonstrates robustness to incomplete observations, with only a moderate decrease in performance.
Our study focuses on small point clouds and a limited number of objects, and we leave broader data modalities and larger-scale settings to future work. Moreover, our method requires fitting a distribution over the prior group, and the choice of the parameterization can impact performance, as seen with the two different prior distributions for $\mathrm{GL}(2)$. Important future directions include improving scalability to high-dimensional or non-compact groups, incorporating additional supervision signals to guide convergence, and integrating discovered symmetries into downstream equivariant networks. This work demonstrates that flow matching offers a promising path toward automatic symmetry discovery, lending us insights into the data and offering improved sample efficiency and generalization for downstream tasks.
Acknowledgements
The authors are grateful to Nima Dehmamy, Minghan Zhu and Rose Yu for helpful discussions. F.E. and J.W.M. would like to acknowledge support from the Bosch Center for Artificial Intelligence. J.W.M. acknowledges additional support from the European Union’s Horizon Framework Programme (Grant agreement ID: 101120237). L.L.S.W. would like to acknowledge support from NSF Grants 2107256. R.W. would like to acknowledge support from NSF Grants 2442658 and 2134178. This work is supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/).
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
LLM Usage
LLMs were used to revise sentences, correct grammar, and draft portions of this work. They were also used to generate parts of the model code and most of the visualization code.
Appendix A Derivation of the Target Flow Field
This section provides a detailed derivation of the target vector field $u_{t}(x_{t}\mid x_{1})$ used in our Lie group flow matching framework. We begin by introducing flow matching on Lie groups, followed by the derivation of the target vector field on the data space induced by the group action.
Interpolants on Lie Groups.
Denote the left action of $G$ on itself for any $g\in G$ as $L_{g}:G\to G,\hat{g}\mapsto g\hat{g}$. Let $T_{g}G$ be the tangent space at $g\in G$. The push-forward of the left action is $(L_{g})_{*}:T_{\hat{g}}G\to T_{g\hat{g}}G$, which transforms tangent vectors at $T_{\hat{g}}G$ to tangent vectors at $g\hat{g}\in G$ via the left action.
To derive FM on a Lie group, we need to define the interpolant as in the Euclidean case. An exponential curve starting at $g_{0},g_{1}\in G$ can be defined by (Sherry and Smets, 2025):
$\gamma:[0,1]\to G;\quad t\mapsto g_{0}\exp(t\log(g_{0}^{-1}g_{1})).$ (7)
This curve starts at $\gamma(0)=g_{0}$ and ends at $\gamma(1)=g_{1}$. The multiplicative difference is $g_{0}^{-1}g_{1}$ and the logarithm maps it to the Lie algebra, giving the direction from $g_{0}$ to $g_{1}$. The scaled version is then exponentiated, recovering intermediate group elements along this curve. Using these exponential curves, we can define the target vector field as
$u_{t}(g_{t}\mid g_{1})=\frac{(L_{g_{t}})_{*}\log(g_{t}^{-1}g_{1})}{1-t}.$ (8)
Defining the Target Vector Field.
Note that we couple source and target by an explicitly sampled group action. According to (4), given $x_{1}\sim q$ and $g\sim p(G)$, and set $x_{0}=gx_{1}$ and $A:=\log(g^{-1})\in\mathfrak{g}$. Consider the group curve $g_{t}=g\exp(tA)$ so that $g_{t}^{-1}=\exp((1-t)A)$ and the induced path $x_{t}=g_{t}x_{1}=x_{0}\exp(tA)$. Plugging $g_{1}=e$ into the group interpolant (8), we have:
$u_{t}(g_{t}\mid e)=\frac{(L_{g_{t}})*\log(g_{t}^{-1})}{1-t}=(L_{g_{t}})*A,$ (9)
which is a time independent target field in the Lie algebra along this curve. To obtain the conditional target on the data path, we first define the orbit map $\phi_{x_{1}}:G\to\mathcal{X},\ \phi_{x_{1}}(g)=gx_{1}$. Its differential at $g_{t}$ is given as $(d\phi_{x_{1}})_{g_{t}}:T_{g_{t}}G\longrightarrow T_{x_{t}}\mathcal{X}$, where $x_{t}=\phi_{x_{1}}(g_{t}).$ Note that $(L_{g_{t}})*:\mathfrak{g}\to T_{g_{t}}G$. Compose them together, we obtain the pushforward operator $J_{x_{t}}:=(d\phi_{x_{1}})_{g_{t}}\circ(L_{g_{t}})*:\mathfrak{g}\longrightarrow T_{x_{t}}\mathcal{X}.$ Then, we push the group target to the data path and obtain the conditional target on data:
$u_{t}(x_{t}\mid x_{1})=(d\phi_{x_{1}})_{g_{t}}\big[u_{t}(g_{t}\mid e)\big]=(d\phi_{x_{1}})_{g_{t}}\big[(L_{g_{t}})*A\big]=J_{x_{t}}(A)\ \in\ T_{x_{t}}\mathcal{X}.$ (10)
We define the stabilizer at $x_{t}$ as $\mathrm{Stab}(x_{t})=\{s\in G:\ s\cdot\!x_{t}=x_{t}\}$ and $\mathfrak{s}_{x_{t}}$ is the Lie algebra of the stabilizer. According to the Orbit-stabilizer theorem, we have:
$\mathrm{Ker}J_{x_{t}}=\mathfrak{s}_{x_{t}},\qquad\mathrm{Im}\,J_{x_{t}}=T_{x_{t}}\!\big(G\!\cdot\!x_{1}\big).$
This means that if the stabilizer of the data is trivial, the pushforward operator is bijective. Consequently, every tangent vector $u_{t}(x_{t}|x_{1})\in T_{x_{t}}(G\cdot x_{1})$ has a unique Lie-algebra coordinate $A\in\mathfrak{g}$ such that $u_{t}(x_{t}\mid x_{1})=J_{x_{t}}(A)$. Using this coordinate identification, for simplicity, we can write $u_{t}(x_{t}|x_{1})\ \equiv\ A$ as the target vector field.
Appendix B Dataset
(a) 2D arrow
(b) 3D irregular tetrahedron
Figure 5: Canonical objects used for generating the 2D (a) and 3D (b) datasets
Figure 6: Samples of the ModelNet10 rotated under $\mathrm{Tet}$ group.
2D arrow.
For the 2D synthetic datasets, we use canonical arrow object as shown in Figure 5a. We generate datasets with discrete symmetries $C_{4}$.
3D irregular tetrahedron.
We use an irregular tetrahedron with no self-symmetries as shown in Figure 5b and generate datasets containing $\mathrm{Tet},\mathrm{Oct},\mathrm{Ico}$ and $\mathrm{SO}(2)$ symmetries.
ModelNet10.
To construct the rotated ModelNet10 dataset, we randomly select 10 objects from each category of ModelNet10 (Wu et al., 2015) as canonical shapes and uniformly sample 128 points from each surface to form point clouds. Each data point is then generated by independently sampling a canonical shape and a transformation from the target group and applying the transformation to the point cloud. Figure 6 shows some samples of the ModelNet10 rotated under $\mathrm{Tet}$ group.
Appendix C Training Details
Sampling Groups from Prior Distribution.
For 2D arrow experiments, the prior known groups are $\mathrm{SO}(2),\mathrm{GL}(2,\mathbb{R})^{+}$. For $\mathrm{SO}(2)$, we use a uniform distribution over the parameter $\theta\sim\mathcal{U}[-\pi,\pi]$ and generate the $2\times 2$ rotation matrices. For the $\mathrm{GL}(2)^{+}$ groups, we use two different sample strategies:
  • Lie algebra coefficients sampling: We sample the coefficients $a,b,c,d\sim\mathcal{U}[-\pi,\pi]$ and construct the Lie algebra element as $A=\begin{bmatrix}a&b\\ c&d\end{bmatrix}$. Then we exponentiate it to obtain the transformation matrix $g=\exp(A)$.
  • Matrix composition based sampling: We sample rotation angle $\theta\sim\mathcal{U}[-\pi,\pi]$, scaling factors $s_{x},s_{y}\sim\mathcal{U}[0.5,2.0]$ and shear factors $s_{h}\sim\mathcal{U}[-1,1]$. Then we construct the transformation matrix as $g=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}s_{x}&0\\ 0&s_{y}\end{bmatrix}\begin{bmatrix}1&s_{h}\\ 0&1\end{bmatrix}$.
Then, we reject any sampled transformation matrices with determinants outside $(0,2]$ to ensure they belong to $\mathrm{GL}(2)^{+}$ and make numerical stable.
For 3D experiments, we consider the prior group $\mathrm{SO}(3)$ and use a uniform distribution over it by using Gaussian normalization over unit quaternions and transforming them into $3\times 3$ matrices.
Training setting for Synthetic data
We use a time-conditioned MLP for flow matching that combines a sinusoidal time embedding module with a 5-layer MLP. The time embedding module uses sinusoidal positional encoding to transform the scalar time $t$ into a hidden dimensional vector, which is then projected to input dimension through a two-layer MLP with ReLU activations. The main network processes the concatenation of the flattened input and the time embedding through four hidden layers and an output layer with ReLU activations. The hidden dimension is set to $128$ for 2D arrow experiment and $512$ for 3D irregular tetrahedron experiment. We use the Adam optimizer with a learning rate of $3\times 10^{-3}$ for both experiments.
Training setting for ModelNet10
For the ModelNet10 experiment,the prediction network adopted is a two-layer Transformer encoder with four attention heads and a hidden dimension of $256$. To condition the network on the time $t$, we use $64$-dimensional sinusoidal time embeddings and concatenate them to the point features. The resulting per-point representations are aggregated via mean pooling to form a global feature. Then, the global feature is then passed through a three-layer MLP to predict a 3D Lie algebra element. AdamW optimizer is used with a learning rate of $1\times 10^{-4}$ and weight decay of $1\times 10^{-2}$.
Time comparison
We conduct training time comparisons between LieFlow and LieGAN on the Modelnet10 datasets. Both methods are trained on a single GeForce RTX 4090 GPU. Fixing batch size to $512$, for one epoch, LieFlow takes approximately $18$ seconds, while LieGAN takes around $50$ seconds.
Appendix D Analysis on Last-Minute Mode Convergence
Given the challenges in learning the icosahedral group, we hypothesize that using flow matching for discovering discrete groups is in fact quite a difficult task. We illustrate with an even simpler scenario of flow matching directly on group elements (matrices), from $\mathrm{SO}(2)$ to $\mathrm{C_{4}}$. As the target vector field $u_{t}$ and the probability path $p_{t}(x_{t}\mid x_{1})$ are important quantities in our FM objective ((5)), we visualize them for a single sample in Figure 7.
Figure 7: $\mathrm{SO}(2)$ to $C_{4}$ over group elements: target vector field and probability paths.
Each target mode $x_{1}$ is colored differently and the velocities (group difference) to each $C_{4}$ group element are shown with the colored arrows. Near $t=0$, the probability path $p_{t}$ is near uniform as $x_{0}$ is essentially noise and can be far away from $x_{1}$, i.e. there is still a lot of time $t$ for it to move to any mode. As $p_{t}$ is nearly uniform, the average of the velocities dominate and pulls $x_{t}$ close to the middle of the red and blue modes. Here, the vector field produces nearly 0 mean velocity, as it is equidistant to the red and blue modes, and equidistant to the green and yellow modes. The distribution $p_{t}$ becomes peakier towards the closer modes (red and blue), but still remains somewhat uniform. As such, the training target averages out close to 0. Near the end of training, when $t$ is close to $1$, here the velocities become larger (due to the scaling factor $t$ inside the exponential) and $p_{t}$ now becomes more uni-modal, picking the closest mode (as there is little time left to go to the other modes, making them more unlikely). Thus, our task of finding a subgroup within a larger group is challenging precisely because the modes are “symmetric” (by definition of the subgroup).
(a) $H_{p}(x_{1}\mid x_{t})$
(b) $p(x_{1}\mid x_{t})$
Figure 8: Left: The entropy of the posterior $p(x_{1}\mid x_{t})$ is generally uniform until $t\approx 1$.
Right: $x_{0}$ tends to converge to the closest mode/component (e.g., $x_{0}$ closest to $\pi$ tends to converge to $\pi$), but the posterior remains nearly uniform until $t\approx 1$.
Flow Matching Directly on Group Elements.
To verify our hypothesis, we perform flow matching directly on the group elements from $\mathrm{SO}(2)$ to $C_{4}$. The group elements can be parameterized by a scalar $\theta$, allowing us to analyze the behavior of the learned vector field more easily. Each data sample $x_{1}$ is a scalar from the set $\{-\pi,-\pi/2,0,\pi/2\}$ and the source distribution is $q=\mathcal{U}[-\pi,\pi]$. We use the same network architecture as in the 2D experiments. We can further compute the reverse conditional or the posterior $p_{t}(x_{1}\mid x_{t})$ that can help verify two things: 1) that $x_{t}$ remain nearly stationary until $t$ approaches $1$, i.e. $p(x_{l}\mid x_{t})$ remains nearly uniform, and 2) $x_{t}$ chooses the closest group element $x_{1}$. The exact details of how $p_{t}(x_{1}\mid x_{t})$ is computed is given in Appendix D.2.
Figures 10 in Appendix D.1 visualizes the progression of $x_{t}$ over time for $100$ samples, showing that $C_{4}$ symmetry is clearly identified. Figures 8(a) plots the entropy of the posterior, and shows that even in this simple scenario, the entropy remains near the maximum (uniform distribution) until $t$ approaches $1$. Figure 8(b) visualizes the posterior $p(x_{1}\mid x_{t})$ averaged over 1,0001{,}000 samples and separated into $4$ classes, depending on the original $x_{0}$ and its nearest mode in $\{-\pi,-\pi/2,0,\pi/2\}$ (only two shown, other classes are similar, see Figures 11(b) in Appendix D.1). For all 4 classes, we observe that $p(x_{1}\mid x_{t})$ remains nearly uniform until around $t=0.9$, when it becomes uni-modal to the closest mode/component, e.g., for all $x_{0}$ that were closest to the mode $\mu=-\pi$, then the posterior $p(x_{1}\mid x_{t})$ converges to $\mu$. This shows both that the learned vector field remains nearly 0 for most timesteps, only converges when little time is left, and converges to the closest group element.
Note that this differs from the standard FM setting, where the flow is defined on a high-dimensional space. Bertrand et al. (2025) find that for high-dimensional flow matching with images, the entropy of the posterior drops at small times, transitioning from a stochastic phase that enables generalization to a non-stochastic phase that matches the target modes. In our low-dimensional setting of symmetry discovery, we find the opposite occurs: the entropy stays near uniform until large $t$, where it suddenly converges. We term this phenomenon “last-minute mode convergence”.
Figure 9: $\mathrm{SO}(2)$ to $C_{4}$ group elements: Evolution of $x_{0}$’s and their velocities over time.
Figure 9 shows the evolution of a uniform grid over $\mathrm{SO}(2)$ over time. As we hypothesized, we can see from the velocity arrows that particles move slowly closer to the midpoint between two modes ($t=0$ to $t=0.60$) but the velocities are relatively small. From $t=0.70$ onwards, we can see that the shift towards the closest modes and the velocities become increasingly large until $t=1$. Figures 12 (Appendix D.1) shows a clearer picture of how the velocity evolves over time. From $t=0$ to $t=0.6$, we see that the $x_{t}$’s are nudged towards the midpoints between modes. Interestingly, we see some oscillating behavior where the vector field flips sign, suggesting that the net velocity hovers close to 0 with no defined pattern. From $t=0.8$ onwards, we see that points $x_{t}$ are pushed to the closest $x_{1}$ mode, with increasing velocities. With near zero net velocity until $t=0.6$, these plots also support the observation of “last-minute mode convergence”.
D.1 Flow Matching on Group Elements
This section contains additional visualizations for the flow matching experiment on group elements from $\mathrm{SO}(2)$ to $C_{4}$ as described in Section D.
Figure 10: $\mathrm{SO}(2)$ to $C_{4}$ group elements: Visualization of $100$ samples over time.
Figures 10 shows the time progression of $x_{t}$ when generating $100$ samples from $t=0$ to $t=1$. We can see that the samples converge to the closest group element in the orbit.
Figures 11(b) in Figures 11 shows the posterior $p(x_{1}\mid x_{t})$ for all $4$ classes, depending on the original $x_{0}$ and its nearest mode in $\{-\pi,-\pi/2,0,\pi/2\}$. We can see that for all $4$ classes, the posterior $p(x_{1}\mid x_{t})$ remains nearly uniform when $t$ closes to $1$.
(a) $H_{p}(x_{1}\mid x_{t})$
(b) $p(x_{1}\mid x_{t})$
Figure 11: Visualizations of the posterior $p(x_{1}\mid x_{t})$. The left graph shows the entropy of the posterior is generally uniform until $t\approx 1$. The right graph visualizes $p(x_{1}\mid x_{t})$ for each target mode and demonstrates that the generated samples, depending on the position of $x_{0}$, converge to the closest mode.
Figures 11 visualizes the posterior $p(x_{1}\mid x_{t})$, divided into 4 buckets depending on the initial position $x_{0}$. We see that the posterior remains nearly uniform until $t=0.95$, and then $x_{t}$ converges to the mode that was the closest when $t=0$.
Figure 12: $\mathrm{SO}(2)$ to $C_{4}$ group elements: velocity phase portraits over time.
Figures 12 visualizes how the velocities change over time. Until $t=0.6$, the velocities are close to 0 but slightly push the $x_{t}$’s towards the midpoints between modes. From $t=0.8$ onwards, we see that points $x_{t}$ are pushed to the closest $x_{1}$ mode, with increasing velocities.
D.2 Derivation of Posterior Computation
To compute the posterior, we use Bayes’ rule as
$p(x_{1}|x_{t})=\frac{p(x_{t}|x_{1})p(x_{1})}{p_{t}(x_{t})},$ (11)
where $p(x_{t}\mid x_{1})$ is the likelihood and $p(x_{1})$ is the data distribution. Since flow matching learns a velocity field rather than explicit conditional distributions, we cannot directly compute the likelihood and need to invert the flow and integrate the continuity equation.
We first sample initial points $x_{0}$ from a uniform prior over $\pi,\pi)$ and forward simulate using the learned velocity field $v_{\theta}$ to obtain the trajectories $x_{t}$ and their log probabilities through the continuity equation. To evaluate the likelihood $p(x_{t}\mid x_{1})$ for each candidate $x_{1}$, we invert the linear interpolation $x_{t}=(1-t)x_{0}+tx_{1}$ to find $x_{0}=(x_{t}-tx_{1})/(1-t)$, then integrate the divergence along the trajectory from this inverted $x_{0}$. Special handling is required at the boundaries: at $t=0$, the posterior equals the prior $p(x_{1})$ (which we know in this simple scenario) due to independence, while near $t=1$, we use a Gaussian approximation to avoid numerical instability from division by (1t)(1-t).
Algorithm 4 Compute Posterior $p(x_{1}\mid x_{t})$ for Flow Matching
0: Velocity field $v_{\theta}(x,t)$, prior $p_{0}\sim\mathcal{U}[-\pi,\pi)$, target $p_{1}(x_{1})=\sum_{k=1}^{K}\frac{1}{K}\delta(x_{1}-\mu^{(k)})$, time steps $T$
1:Initialize: Sample $\{x_{0}\}^{N}\sim p_{0}$
2:
3:Forward Simulation:
4:for $t\in\{0,\frac{1}{T},\ldots,1\}$ do
5:  $x_{t}\leftarrow\text{ODESolve}(v_{\theta},x_{0},[0,t])$ $\triangleright$ Integrate velocity field
6:  $\log p_{t}(x_{t})\leftarrow\log p_{0}(x_{0})-\int_{0}^{t}\nabla\cdot v_{\theta}(x_{s},s)ds$ $\triangleright$ Continuity equation
7:end for
8:
9:Posterior Computation:
10:for $t\in\{0,\frac{1}{T},\ldots,1\}$ do
11:  if $t=0$ then
12:   $p(x_{1}\mid x_{0})\leftarrow p_{1}(x_{1})$ $\triangleright$ Independent at $t=0$
13:  else if $t\approx 1$ then
14:   for $k=1$ to $K$ do
15:    $\log p(x_{t}\mid x_{1}^{(k)})\leftarrow-\frac{\|x_{t}-x_{1}^{(k)}\|^{2}}{2\sigma^{2}}$, where $\sigma=0.1$ $\triangleright$ Gaussian Approximation
16:   end for
17:  else
18:   for $k=1$ to $K$ do
19:    $x_{0}^{(k)}\leftarrow\frac{x_{t}-t\cdot x_{1}^{(k)}}{1-t}$ $\triangleright$ Invert linear interpolation
20:    Forward simulate path from $x_{0}^{(k)}$ to time $t$ using $v_{\theta}$ and continuity equation
21:    $\log p(x_{t}\mid x_{1}^{(k)})\leftarrow\log p_{0}(x_{0}^{(k)})-\int_{0}^{t}\nabla\cdot v_{\theta}(x_{s},s)ds$
22:   end for
23:  end if
24:  for $k=1$ to $K$ do
25:   $\log p(x_{1}^{(k)}\mid x_{t})\leftarrow\log p(x_{t}\mid x_{1}^{(k)})+\log p_{1}(x_{1}^{(k)})$ $\triangleright$ Bayes’ Rule
26:  end for
27:
28:  $p(x_{1}^{(k)}\mid x_{t})\leftarrow\frac{\exp(\log p(x_{1}^{(k)}\mid x_{t}))}{\sum_{k}\exp(\log p(x_{1}^{(k)}\mid x_{t}))}$ $\triangleright$ Normalize over $K$ components
29:end for
30:
31:Return Posterior $\{p(x_{1}\mid x_{t})\}$ for all time steps
Appendix E Power time schedule
Figure 13 shows the density of the power distribution for different skewness values.
Figure 13: Power distribution density
Figures 14 compares the time progression of $x_{t}$ between without and with power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ dataset. Figure 14(a) shows that without power time schedule, the samples remain nearly stationary until $t$ approaches $1$, while in Figure 14(b), with power time schedule, the samples start moving towards the target modes earlier, even at small $t$ values. This indicates that the power time schedule helps to alleviate the last-minute mode convergence issue discussed in Section 5.4 and Appendix D.
(a) Time progression of $x_{t}$ without power time schedule
(b) Time progression of $x_{t}$ with power time schedule
Figure 14: Test progression comparison between without and with power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ dataset.
E.1 Ablation study on skewness
We perform an ablation study on the skewness parameter of the power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ setting. Table 4 reports the Wasserstein-1 distance between the generated transformations and the ground truth group elements for different skewness values. Figure 15 visualizes the training curves for different skewness values. From the results, we observe that increasing the skewness of the power time schedule leads to better performance and faster convergence. Since the improvement saturates around skewness $5$, we use this value for all 3D experiments in the main paper.
Table 4: Ablation study on skewness of power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ setting. Wasserstein-1 distance between the generated transformations and the ground truth group elements are reported. Lower is better.
skewness 1 (uniform) 2 3 4 5
$0.328$ $0.171$ $0.137$ $0.121$ $0.118$
Figure 15: Ablation study on the skewness of power time schedule for $\mathrm{SO}(3)$ to $\mathrm{Ico}$ setting. $y$ axis denote the Wasserstein-1 distance between the generated transformations and the ground truth group elements, and $x$ axis denote training epoch.
Appendix F Additional Results
This section contains additional visualization results
F.1 2D arrow
Figure 16: $\mathrm{GL}(2)^{+}$ to $C_{4}$: Time progression of $x_{t}$ when generating $5$ samples over 2020 steps. The gray arrow shows the original $x_{1}$ and the color represents the determinant of the generated transformation matrix.
Figures 16 shows the time progression of $x_{t}$ when generating $5$ samples over $20$ steps for $\mathrm{GL}(2)^{+}$ to $C_{4}$. The gray arrow shows the original $x_{1}$ and the color represents the determinant of the generated transformation matrix. We can see that the transformed point clouds converge to the closest group element in the orbit.
F.2 3D irregular tetrahedron
(a) $\mathrm{SO}(3)$ to $\mathrm{Tet}$
(b) $\mathrm{SO}(3)$ to $\mathrm{Oct}$
(c) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$
(d) $\mathrm{SO}(3)$ to $\mathrm{Ico}$ (with power time schedule)
Figure 17: A 2D PCA visualization of trajectories from t = 0 (left) to t = 1 (right) of the centroids of the transformed objects over 100 samples.
Figures 17 visualizes the trajectories of the centroids over time for $100$ samples, and PCA is performed to project them onto the 2D plane. We can see that our model seems to learn $\mathrm{Tet}$, $\mathrm{Oct}$ $\mathrm{Ico}$ and $\mathrm{SO}(2)$ elements, creating some visible clusters.
Figure 18: $\mathrm{SO}(3)$ to $\mathrm{Tet}$: time progression of $x_{t}$ when generating $5$ samples over $100$ steps. The transparent tetrahedron indicates the original $x_{1}$.
Figure 19: $\mathrm{SO}(3)$ to $\mathrm{Oct}$: time progression of $x_{t}$ when generating $5$ samples over $100$ steps. The transparent tetrahedron indicates the original $x_{1}$.
Figure 20: $\mathrm{SO}(3)$ to $\mathrm{Ico}$: time progression of $x_{t}$ when generating $5$ samples over $100$ steps. The transparent tetrahedron indicates the original $x_{1}$. Training with power time schedule.
Figure 21: $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$: time progression of $x_{t}$ when generating $5$ samples over $100$ steps. The transparent tetrahedron indicates the original $x_{1}$.
We also show the intermediate $x_{t}$ during generation in Figure 18, Figure 19, Figure 20 and Figure 21 for $\mathrm{Tet}$, $\mathrm{Oct}$, $\mathrm{Ico}$ and $\mathrm{SO}(2)$ respectively. For the $\mathrm{Tet}$, $\mathrm{Oct}$, and $\mathrm{Ico}$, we see that the transformed point clouds converge to the closest group element in the orbit instead of the original $x_{1}$, which shows that the model has learned the symmetry structure of the data. For the $\mathrm{SO}(2)$ group, we can see that the model produces rotations to turn one node of the tetrahedron to be aligned with the $z$-axis and one triangle face to be in the $xy$-plane, which corresponds to $\mathrm{SO}(2)$ symmetries around the $z$-axis.
F.3 ModelNet10
Visualization results for rotated ModelNet10 are shown in Figures 22. The generated group elements are clustered around the ground truth group elements for all target groups, which means that our model successfully discovered the symmetries in the rotated ModelNet10 dataset. Note that for the $\mathrm{Ico}$ group, the situation here is even worse than the irregular tetrahedron case without power time schedule, as the model fails to discover any symmetry due to the increased complexity of the dataset. However, with the power time schedule, the model still is able to discover the $\mathrm{Ico}$ symmetries, even with some points masked during training. This shows the effectiveness and robustness of our model combined with the power time schedule in discovering complex symmetries in real-world datasets.
(a) $\mathrm{SO}(3)$ to $\mathrm{Tet}$
(b) $\mathrm{SO}(3)$ to $\mathrm{Oct}$
(c) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over z axis
(d) $\mathrm{SO}(3)$ to $\mathrm{SO}(2)$ over $(0,\frac{1}{2},-\frac{\sqrt{3}}{2})$ axis
(e) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, without power schedule
(f) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule
(g) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule, 5% masked points
(h) $\mathrm{SO}(3)$ to $\mathrm{Ico}$, with power schedule, 10% masked points
Figure 22: Modelnet10 results: Visualization of 5,0005{,}000 generated elements of $\mathrm{SO}(3)$ by converting them to Euler angles. The first two angles are represented spatially on the sphere using Mollweide projection and the color represents the third angle. The elements are canonicalized by the original random transformation and the ground truth elements of the target group are shown in circles with gray borders. The points are jittered with uniformly random noise to prevent overlapping.